OpenStudy (basketball12345):

Find the standard form of the equation of the parabola with a focus at (3, 0) and a directrix at x = -3.

OpenStudy (adonine):

any Ideas?

OpenStudy (welshfella):

this will be a parabola passing through the origin and opening to the right

OpenStudy (welshfella):

The general form of these is y^2 = 4ax where the focus is at (a,0)

OpenStudy (welshfella):

in this case what is the value of a?

OpenStudy (welshfella):

a is the x cordinate of the focus

OpenStudy (basketball12345):

so a is 3?

OpenStudy (basketball12345):

would the answer be y^2=6x? @welshfella

OpenStudy (welshfella):

yes a is 2 so the answer is y^2 = 4*3a